two men startedtogether to walk a distance at 3 km per hour and 4 km per hour respectivelythe former arrives 30 mts before the latter find distance the walked

Answer by Alan3354(69443)

two men startedtogether to walk a distance at 3 km per hour and 4 km per hour respectivelythe former arrives 30 mts before the latter find distance the walked
--------
Pitiful typing.
----
I'll assume 30 mts means 30 minutes = 0.5 hours
---
d = r*t
r = slower rate
3*(t+0.5) = 4t
3t + 1.5 = 4t
t = 1.5 hours
----
d = 4t = 6 km
-----------
PS the "respectively" serves no purpose, tho that was probably in the question you copied.